40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 14002     Entry time: Fri Jun 22 00:06:13 2018
Author: shruti 
Type: Update 
Category: General 
Subject: over-head fluorescent lights down 

Two out of the four over-head fluorescent lights in the X end of the interferometer were flickering today.

ELOG V3.1.3-